A 0.150-kg frame, when suspended from a coil spring, stretches the spring 0.050 m. A

0.200-kg lump of putty is dropped from rest onto the frame from a height of 30.0 cm (Fig.).

Find the maximum distance the frame moves downward from its initial position.

the

spring to

find the force constant

k

of the spring. Let

s

be the amount the spring is stretched.



1
Expert's answer
2022-05-06T15:34:17-0400

"mgh=\frac{Mv^2}2+\frac{kx^2}2,"

"h=\frac{Mv^2+kx^2}{2mg},"

"h=15~cm."